OpenStudy (anonymous):

Medal and Fan Rewarded!! A juice company did a survey among 100 customers to find their juice preferences. The customers were asked about their preferences for apple or orange juices. Out of the total 35 people who liked apple juice, eight liked orange juice also. There were 20 people who liked orange juice. Part A: Summarize the data by writing the values that the alphabets A to I in the table below represent. (5 points) Part B: What percentage of the total people did not like either apple or orange juice Part C: Do the people surveyed have more dislike for apple juice than orange juice? Justify your answer. (2 points)

OpenStudy (anonymous):

@TuringTest @tester97 @Compassionate @kirbykirby

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ayoubEpst

OpenStudy (kirbykirby):

I don't see the table?

OpenStudy (anonymous):

oh sorry ill draw it for you real quick (:

OpenStudy (anonymous):

 sorry that i took so long

OpenStudy (anonymous):

@kirbykirby you there??

OpenStudy (kirbykirby):

oops sorry I was answering someone else

OpenStudy (anonymous):

its ok ^-^

OpenStudy (kirbykirby):

If 35 liked Apple, then 100-35 = 65 didn't like apple juice So, you know that G = 35 (the total that like apple juice) and H = 65 (the total that don't like apple juice). ANd the total number of people is 100, so I = 100. A = 8, since this is the number of people that like Apple and Orange juice (given in your text) So, those that like Apple but not orange juice is 35 - 8 = 27, so D=27 You know that the number of people who like orange juice is 20, so C = 20 But since A + B = C, that means B = C - A B= 20 - 8 B= 12 Now, since C + F = I Then: F = I - C F= 100 - 20 F=80 Since D + E = F then, E = F - D E = 80 - 27 E = 53

OpenStudy (anonymous):

oh ok got it so how would go with it having to a,b,c ect jw

OpenStudy (kirbykirby):

well this^ above answer part a). Are you asking how to do part b) and c) ?

OpenStudy (anonymous):

well i was just wondering how i would summarize the data by writting the values in alphabets like from a to i if its no trouble..

OpenStudy (kirbykirby):

All the letters I wrote up there, like A = 8, B = 12, etc. just put those in the table corresponding to the right letter

OpenStudy (anonymous):

ohh ok im so sorry i totally didnt see that im really sorry. but you dont mind helping with part and C do u ??

OpenStudy (kirbykirby):

They want you to compare how many people dislike orange juice vs. how many people dislike apple juice. The total number of people who dislike orange juice is represented by F = 80 The total number of people who dislike apple juice is represented by H = 65 So you see that more people dislike orange juice than apple juice. (Your question asks do people have more dislike for apple juice than orange? The answer is No, because less people dislike apple compared to orange)

OpenStudy (anonymous):

ok got it (: so than Part b is asking for a percentage and i got 7.2 is that correct?

OpenStudy (anonymous):

im confused on how to figure out the percentage for part c

OpenStudy (anonymous):

@kirbykirby you there??

OpenStudy (kirbykirby):

oye sorry :S There is no percentage to calculate for part c). For part b): Say \(A\) = did not like apple \(O\) = did not like orange, Then \(P(A ~or~O)=P(A \cup O)\\=P(A)+P(O)-P(A \cap O)\\=\dfrac{65}{100}+\dfrac{80}{100}-\dfrac{53}{100}=\dfrac{92}{100}=92\% \)

OpenStudy (kirbykirby):

Hmm maybe that is too complicated for this level. An easier way to look at part b) is to see that those who liked Apple and Orange was 8 people out of 100. So the number that don't like apple or orange is 100 - 8 = 92. Since you want a percentage, 92/100 = 92%

OpenStudy (anonymous):

oh ok thank you so so much im really sorry for all this iv been on this test since 6am :S and i got 1 question left

OpenStudy (anonymous):

i have 1 question how did u get 8 ??

OpenStudy (kirbykirby):

That was given in your problem. You know that out of 35 people who like apple juice, 8 liked orange. Meaning that 8 people like both apple and orange juice

OpenStudy (anonymous):

ohh ok sorry i see now (; thx

OpenStudy (kirbykirby):

=]

OpenStudy (anonymous):

i have another question lol sorry for part a im confused on what letter b represents?

OpenStudy (kirbykirby):

B is those that Like orange juice, but don't like apple juice. I got the result of B knowing that A + B = C (C is the total of those that like orange juice.. it is the row total)

OpenStudy (kirbykirby):

C + F = I because I is the column total. So to find F, by algebra: subtract C from both sides: F = I - C Substitute the values you know for I and C: F= 100 - 20 F=80

OpenStudy (kirbykirby):

yep

OpenStudy (kirbykirby):

yes

OpenStudy (anonymous):

im really sorry for keeping u up i would seriously give like a thousand s of medals if it was possible not even joking lol :P

OpenStudy (kirbykirby):

loll oh don't worry about it :)